OpenStudy (anonymous):

The city of Port St.Lucie is considered the second fastest growing city in Florida. In June of 2002, it had a reported population of 9.7x10^4 people. By, June of 2004 the city's population had reached 1.18x10^5 people. How much did Port St.Lucie's population increase, in thousands, from June 2002 to 2004 ?

OpenStudy (rushwr):

\[1.8 * 10^{5}= 18 * 10^{4}\] So now subtract 9.7 * 10^4 from 18 * 10^4

OpenStudy (rushwr):

\[18*10^{4} - 9.7 * 10^{4} = 8.3 * 10^{4}\]

OpenStudy (rushwr):

So now write it in thousands !

OpenStudy (anonymous):

I thought it was \[9.7x10^4 - 18x10^5\]

OpenStudy (rushwr):

oh no that's not it ! As we can see in the question there's an increment in the population isn't it ? So we have to subtract the smaller value from the big value ! So when they ask you to find the increment they are asking you to find the difference between the population in 2002 and 2004 !

OpenStudy (anonymous):

Ohhh I get it now THANK YOU sooo much !!!!!!!

OpenStudy (rushwr):

No problem ! As long as you get it !

OpenStudy (anonymous):

So, the population increase would be 8,300 ?

OpenStudy (rushwr):

8.3 * 10^4 = 83 000

OpenStudy (rushwr):

You have to move 4 decimals to the right ! Isn't it ?

OpenStudy (rushwr):